Mary deposited $5500 in an account earnings 5.5 % interest. How long will it take for her investment to earn $205 in interest? Round to the nearest day.
I would like to know how to solve this equation.
Thank you!

Found 2 solutions by ewatrrr, stanbon:
Answer by ewatrrr(24785)

Important to know if it is simple or compounding interest

Answer by stanbon(75887)

Mary deposited $5500 in an account earnings 5.5 % interest. How long will it take for her investment to earn $205 in interest? Round to the nearest day.
--------
Interest = P*r*t
205 = 5500*0.055*t
-----
t = 0.6777 years
-----
t = 0.6777*365 = 247.35 days
---------------------------------
Cheers,
Stan H.
-----------------